The graphs of each group of equations have at least one characteristic in common. Name the
characteristic(s)
2x-4y=12, 3y+2x=-2, x+y=0

Answer by stanbon(75887)

The graphs of each group of equations have at least one characteristic in common. Name the
characteristic(s)
2x-4y=12
slope = 2/4 = 1/2
int = -3
---------------
3y+2x=-2
slope = -2/3
int = -1
-----------------
x+y=0
slope = -1
int = 0
-----
Ans: Each graph is a line.
-----------
Cheers,
Stan H.
-----------